• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Perdrisat

    Log entry time 00:56:45 on March11,2006

    Entry number 166204

    keyword=swing shift journal

    various problems in the hall are being fixed (amplifiers for the veto counters in ND overload NIM crate; HV for drift chamber redistributed).
    Frist production run # 2849 starts at 19:13; checking current in chambers. Accidentals 25%.
    HW plate put in at 20:00
    NMR being done simultaneously
    snmall changes of HV on drift chambers
    run # 2850 started at 20:03, after change of half-wave plate to IN
    run # 2851 started at 20:48; target oven temperature is dropping since 20:50 (sofar to 200C)
    Aidan replaced controller. Oven temperature bottomed out at 193 C and started rising after 21:17. He claims that loss of 3-He polarization should have been small.

    Submitted late by Bryan Moffit. Hope this was intended to be submitted.